Use the distance formula and the converse of the Pathagorean [sic] Theorem to determine whether the points are vertices of a right triangls. (please show all work)
1.(1.-4),(5,6),(-2,3)
2.(-2,4),(5,3),(0,-1)
3.(2,1),(6,-3),(-5,1)
THANKS

Answer by Alan3354(69443)

Use the distance formula and the converse of the Pathagorean Theorem to determine whether the points are vertices of a right triangls.(please show all work)
1.(1.-4),(5,6),(-2,3)
2.(-2,4),(5,3),(0,-1)
3.(2,1),(6,-3),(-5,1)
----------------
Label the points
L(1,-4), M(5,6), N(-2,3)
--
Find the length of LM, LN and MN
If the square of the longest = the sum of the squares of the other 2, it's a right triangle. (Unless they're colinear).
-----
LM^2 = diffy^2 + diffx^2 = 4^2 + 10^2 = 116
LN^2 = 3^2 + 7^2 = 58
MN^2 = 7^2 + 9^2 = 130
---
130 <> 110 + 58 --> not a right triangle.
==============
Do the others the same way.